OpenStudy (anonymous):

PLEASSE HELP W THIS PROBLEM: Andy has some $10 bills and some $20 bills in a shoebox under his bed. He has a total of 47 bills worth a total of $610. How many $20 bills does he have?

OpenStudy (anonymous):

a = $10 bills b = $20 bills a + b = 47 -->a = 47 - b 10a + 20b = 610 now sub 47 - b in for a in the 2nd equation 10(47 - b) + 20b = 610 470 - 10b + 20b = 610 -10b + 20b = 610 - 470 10b = 140 b = 14 now sub 14 in for b in either of the equations to find a a + b = 47 a + 14 = 47 a = 47 - 14 a = 33 check... 10a + 20b = 610 10(33) + 20(14) = 610 330 + 280 = 610 610 = 610 (correct) I know you only asked for the number of $20 bills, which is 14, but I had to finish the problem to be able to check my answer.